OpenStudy (campayne):

Complete the square for the expression. Write the resulting expression as a binomial squared. x^2 + 3x + _____

OpenStudy (bonnieisflash1.0):

can't help you anymore

OpenStudy (campayne):

whys that

OpenStudy (3mar):

If I may help?

OpenStudy (campayne):

oh absolutely

OpenStudy (3mar):

Thanks

OpenStudy (3mar):

\[x^2+3x+(\frac{ 3 }{ 2 })^2-(\frac{ 3 }{ 2 })^2\]

OpenStudy (campayne):

yeah i got that exact thing a while ago

OpenStudy (3mar):

then \[(x+\frac{ 3 }{ 2 })^2-\frac{ 4 }{ 9 }\]

OpenStudy (campayne):

yep

OpenStudy (3mar):

Satisfied?

OpenStudy (campayne):

indeed